HUBBLE SPACE TELESCOPE DAILY REPORT #1595 PERIOD COVERED: 0000Z (UTC) 03/20/96 - 0000Z (UTC) 03/21/96 Daily Status Report as of 081/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6359 (Core Properties of the Bulges of Spiral = Galaxies) The WF/PC-2 was used to make snapshot survey observations of = spiral galaxy NGC3885 in order to study its central bulge. The aim is to = investigate the core properties, correlations with global galactic = properties, and the frequency of occurrence of multiple and/or = miscentered nuclei. The observations were completed as scheduled, and no = anomalies were noted. 1.2 Completed FOS/RD 6089 (Eclipse Mapping and Tomography of the UV = Emitting Region of DQ Her) The Faint Object Spectrograph (Red Side) was used to make Rapid = Mode ultraviolet observations of the eclipsing binary star 14Aur-C in = order to do eclipse mapping and tomographic studies. Observations were = made of both the ultraviolet continuum and emission lines. A wavelength = calibration image was also taken. The observations were executed as = scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 5996 (The Morphology of Field Galaxies at High = Redshifts) The WF/PC-2 was used to make observations of several high = redshift field galaxies (Field-1000) in order to study their = morphologies. The observations were completed as planned, and no = anomalies were reported. 1.4 Completed WF/PC-2 6276 (Peculiar and Interacting Galaxies (WC01): = Cycle 5) The WF/PC-2 was used to make observations of galaxy NGC-6745. The = spatial resolution afforded by the HST is expected to reveal important = facts concerning the nature of this object. The observations were = executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Operations Notes: HST operations transitioned to the SAMS back-up facility from = 080/1500Z to 080/2330Z while maintenance was performed on the LAN in the = MOR area. No problems were encountered. The first portion of the ESTR re-conditioning was completed at = 080/1644Z. Another session is to occur today. There was a non-recoverable 12-minute, 11-second realtime data = loss during STGT operations. A TTR was written. As described in HSTAR 5538, during a Solar Array slew from = approximately 68 degrees to 12 degrees at 077/205547Z, the plus wing posit= ion error exceeded the PRS limit of 6 degrees by 0.103 degrees. The = position came back into bounds at 077/205611Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH